A 16-year annuity pays $1,300 per month, and payments are made at the end of each month. The interest rate is 13 percent compounded monthly for the first six years and 12 percent compounded monthly thereafter. 


What is the present value of the annuity?
Answer #2

SOLUTION :


r = 13/12 %  = 13/1200 /per month for first 6 years (n = 72 months).

=> (1 +r) = 1 + 13/1200 = 1213/1200 

A = monthly payment = 1300 ($) 


PV of annuity of first 7 years 

= A((1+r)^n - 1) / (r*(1+r)^n) 

= 1300((1213/1200)^72 - 1) / ((13/1200)*(1213/1200)^72))

= 64760.05 ($)


After 6 years :


r = 12/12 % = 1% = 0.01 per month

=> (1 + r) = 1.01

n = (16 - 6) = 10 years = 120 months.

A = 1300 ($)


PV (at 6 years end) 

= A((1+r)^n - 1) / (r*(1+r)^n) 

= 1300(1.01^120 - 1) / (0.01 * 1.01^120))

= 90610.68  ($)


So, PV at the starting point

= 90610.68 / (1.01)^120

= 27454.56 ($)


So, total PV at the beginning

= 64760.05 + 27454.56

= 92214.61 ($) (ANSWER).

answered by: Tulsiram Garg

> Please correct as under :
So, PV at the starting point
= 90610.68 / (1.01)^72
= 44262.96 ;
So, total PV at the beginning
= 64760.05 + 44262.96
=109023.01 ($) (ANSWER).

Tulsiram Garg Thu, Jan 27, 2022 7:21 AM

> Please correct 4th line as under :

" PV of annuity of first 6 years"

Tulsiram Garg Sat, Jan 29, 2022 7:18 AM
Answer #1
PVOrdinary Annuity = C*[(1-(1+i/(f*100))^(-n*f))/(i/(f*100))]
C = Cash flow per period
i = interest rate
n = number of years I f = frequency of payment
PV= 1300*((1-(1+ 13/1200)^(-6*12))/(13/1200))
PV = 64760.05
Using Calculator: press buttons "2ND"+"FV" then assign
PMT =1300
I/Y =13/12
N =6*12
FV = 0
CPT PV
Using Excel
=PV(rate,nper,pmt,FV,type)
=PV(13/(12*100),12*6,,PV,)
PVOrdinary Annuity = C*[(1-(1+i/(f*100))^(-n*f))/(i/(f*100))]
C = Cash flow per period
i = interest rate
n = number of years I f = frequency of payment
PV= 1300*((1-(1+ 12/1200)^(-10*12))/(12/1200))
PV = 90610.68
Using Calculator: press buttons "2ND"+"FV" then assign
PMT =1300
I/Y =12/12
N =10*12
FV = 0
CPT PV
Using Excel
=PV(rate,nper,pmt,FV,type)
=PV(12/(12*100),12*10,,PV,)
EAR = [(1 +stated rate/no. of compounding periods) ^no. of compounding periods - 1]* 100
? = ((1+13/(12*100))^12-1)*100
Effective Annual Rate% = 13.8032
Future value = present value*(1+ rate)^time
90610.68 = Present value*(1+0.138032)^6
Present value = 41711.19
Using Calculator: press buttons "2ND"+"FV" then assign
FV =-90610.68
I/Y =13.8032
N =6
PMT = 0
CPT PV
Using Excel
=PV(rate,nper,pmt,FV,type)
=PV(0.138032,6,,90610.68,)

Total =41711.19+64760.05

=

106471.24